OpenStudy (anonymous):

The graph of the function is a translation of the graph of f(x) = x2. Graph the function. g(x) = (x − 7)2

OpenStudy (anonymous):

f(x) =3x+2 , g(x) =x-3 g(x) is: 3x+2-x+3 =2x+5

OpenStudy (anonymous):

Find the vertical asymptotes, if any, of the rational function. Do not graph the function. (Enter your answers as a comma-separated list of equations. If an answer does not exist, enter DNE.) f(x) = 7x/(7x + 8)